OpenStudy (anonymous):

Solving inequalities 2x − 1 ≥ (2/3) + (5x/2)

OpenStudy (anonymous):

i have problems with the fractions

OpenStudy (anonymous):

if 2x-1 >= 2/3+5x/2 then 2x-5x/2>=2/3+1 -0.5x>=1.66 -x>=3.33 or x<=3.33

OpenStudy (anonymous):

thank you! i wanna ask you.. why are you soo good at math!

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

lol, thank you. I am a geek :D

OpenStudy (anonymous):

i wish i can be a geek and good at math

OpenStudy (anonymous):

all you need to do is think logically and break down the problem into smaller bits. good luck!

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

wait how you get -0.5x

OpenStudy (anonymous):

5x/2 = 2.5 x 2- 2.5x = -0.5x

OpenStudy (anonymous):

got it thanks